Thank you very much!

Two parallel rails having negligible resistance ar
Answer #1

induced emf in the moving rod = B*v*L

= 0.0118*2*0.1

= 2.36*10^-3 volts

net resistance of the ckt, Rnet = 13 + 6*17/(6+17)

= 17.43 ohms

current through 13 ohm resistor, I = induced emf/Rnet

= 2.36*10^-3/17.43

= 0.135 mA

current through 6 ohms = I*17/(6+17)

= 0.1 mA <<<<<<--------Answer